An ideal solution is made from 6 moles of benzene. & 2.50 Moles off toulene calculate ΔmixG at 298K & 1bar pressure.

Grade:12th pass

1 Answers

Susmriti
5 years ago
∂G=nRT(x1lnx1 + x2lnx2)
Given:-
n=6+2.5= 8.5
R=8.314
T=298K
x1=6/8.5 & x2=2.5/8.5
     =0.7058.    =0.2941
∂G= 8.5×8.314×298×2.303(0.7058log0.7058+ 0.2941log0.2941
∂G= 12757.9 J